OpenStudy (anonymous):

Find eccentricity and directrix r=5/(-1+2cosq)

OpenStudy (experimentx):

The standard equation of the conic is given by \[ \frac{l}{r} = 1 + e\cos\theta\]

OpenStudy (experimentx):

eccentricity = 2

OpenStudy (anonymous):

directrix is -5/2 or 5/2 ??

OpenStudy (experimentx):

are you trying to get directrix in polar form or in xy form??

OpenStudy (anonymous):

directrix is x=-5/2 or 5/2 ?

OpenStudy (experimentx):

must be 5/2 http://www.wolframalpha.com/input/?i=polar+plot+r+%3D5%2F%28-1%2B2cos \theta%29

OpenStudy (anonymous):

THX

OpenStudy (experimentx):

wait, i don't think it should be 5/2

OpenStudy (anonymous):

i think ed = 5 then e=2 so d = 5/2

OpenStudy (anonymous):

center is 10/3

OpenStudy (experimentx):

NO ... the equation of the directrix should be ... something like \[ \frac{\sqrt{r^2 - (5/2)^2}}{r} = \cos \theta\] or \[ \frac{(5/2)}{r} = \sin \theta\]

OpenStudy (experimentx):

yeah ... centre is 10/3

OpenStudy (anonymous):

i am using polar equation r=ed/1-ecosq d=directrix

OpenStudy (experimentx):

Ooops ... sorry!! i was wrong!!

OpenStudy (anonymous):

no nevermind thanks for your help anyway

OpenStudy (experimentx):

but this is hyperbola ... shouldn't the directrix pass though the centre??

OpenStudy (anonymous):

there are two directrix 5/2 and 25/6

OpenStudy (anonymous):

center is in between of two directrix

OpenStudy (experimentx):

I guess it is then ..

OpenStudy (anonymous):

i found a video http://www.youtube.com/watch?v=zV6buIBpuJo